OpenStudy (anonymous):

If 3x^2 + y^2 = 7 then evaluate the second derivative of y with respect to x when x = 1 and y = 2. Round your answer to 2 decimal places. Use the hyphen symbol, -, for negative values.

OpenStudy (anonymous):

On this, I got the first derivative rather easily, but am having difficulty finding the second derivative.

OpenStudy (anonymous):

The first derivative being -3x/y

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

y'=-3x/y is what I meant.

OpenStudy (pawanyadav):

Try for second

OpenStudy (anonymous):

I did try for the second, I just get stuck in a mess.

OpenStudy (pawanyadav):

Its a u/v type function you know how to differentiate them

OpenStudy (pawanyadav):

Or simply 1st = 6x+2y(dy/dx)=0

jimthompson5910 (jim_thompson5910):

\[\Large \frac{dy}{dx} = -\frac{3x}{y}\] is the correct first derivative

OpenStudy (anonymous):

Here's my work for the second derivative, and tell me where I went wrong

OpenStudy (anonymous):

Isn't that what I have?

OpenStudy (pawanyadav):

So 2nd= 6 + 2(dy/dx)^2 +2y (y'')

jimthompson5910 (jim_thompson5910):

yes @Tahamohammed I was confirming you had the right answer

jimthompson5910 (jim_thompson5910):

\[\Large \frac{dy}{dx} = -\frac{3x}{y}\] \[\Large \frac{d}{dx}\left[\frac{dy}{dx}\right] = \frac{d}{dx}\left[-\frac{3x}{y}\right]\] use the quotient rule

OpenStudy (anonymous):

6x+2yy' 6+2y'^2+2yy" y"=6+2y'^2/-2y y"=(6+2(-3x/y)^2)/-2y y"=(6+18x^2/y^2)/-2y y"=(6y^2/y^2+18x^2/y^2)/-2y

OpenStudy (pawanyadav):

Where dy/DX=-3(1)/2=-3/2 So 6+2(-3/2)^2+4(y'')=0 You have to find y''

OpenStudy (anonymous):

Did I make any mistakes so far? @Pawanyadav @jim_thompson5910 ?

OpenStudy (pawanyadav):

Now just plug values you are correct

OpenStudy (anonymous):

Just plug in the values?

OpenStudy (anonymous):

Shouldn't I simplify further?

OpenStudy (pawanyadav):

No need of it ......but if you want to then you can do

OpenStudy (anonymous):

So is the answer -2.625 then?

OpenStudy (anonymous):

-2.63 rounding to the second decimal place.

OpenStudy (pawanyadav):

It round's to -2.62

OpenStudy (pawanyadav):

Done

OpenStudy (anonymous):

Okay, thanks!

OpenStudy (pawanyadav):

You are welcome

OpenStudy (anonymous):

Wait, why is it -2.62? @Pawanyadav

OpenStudy (anonymous):

@Pawanyadav

OpenStudy (pawanyadav):

-2.625 if we get 5 after even number We don't increase it by one Here 5 is after 2.

OpenStudy (anonymous):

I never heard of that rule, I always remembered greater than 5 it goes up 1.

OpenStudy (pawanyadav):

Thats for Greater than 5 not 5

OpenStudy (anonymous):

I meant including 5.

jimthompson5910 (jim_thompson5910):

you should find that \[\Large \frac{d^2y}{dx^2} = \frac{-3y^2 - 9x^2}{y^3}\] Now plug in (x,y) = (1,2) \[\Large \frac{d^2y}{dx^2} = \frac{-3y^2 - 9x^2}{y^3}\] \[\Large \frac{d^2y}{dx^2} = \frac{-3(2)^2 - 9(1)^2}{(2)^3}\] \[\Large \frac{d^2y}{dx^2} = -\frac{21}{8}\] \[\Large \frac{d^2y}{dx^2} = -2.625\] \[\Large \frac{d^2y}{dx^2} = -2.63\]

OpenStudy (anonymous):

So it is -2.63? Thanks everyone!

OpenStudy (pawanyadav):

That's solve...try to find that rule

OpenStudy (anonymous):

I put the answer -2.63 and it said I was right, so I think what I thought was right.

OpenStudy (jmark):

3x^2 + y^2 = 7, differentiating w.r.t x 6x + 2y.y' = 0 or y' = -3x/y differentiating again, 6 + 2y' + 2yy'' = 0 y' = -3x/y = -3/2 putting values, y = 2 and x = 1 y'' = -3/4 = -0.75